Simon walked at 50m/min and then rode his bike at 200m/min for the rest of the way.If it took 33 minutes to complete the whole journey of 3 km, for how long did he walk?
Answer by richwmiller(17219)

W+B=3000,
50*w=W,
200b=B,
We have more info than we need.
We only need
w+b=33,
50*w+200b=3000,
b=9 min
w=24 min